Evaluate the iterated integral by converting to polar coordinates

  1. Evaluate the iterated integral by converting to polar coordinates

xstT89dIpulAAAAAElFTkSuQmCC
Answer #1

2 14-312 (say) I = 1 (0(² + y2) dy dx limits - y2 +00²=4 y zo X=-2 to to y = Ju-x² X=2 y = Sunxa 30-2 I limit in polar form 8

Doubts are welcome.

Thank You !